Solve:
2x(x+5)=100

Answer by stanbon(75887)

Solve:
2x(x+5)=100
x(x+5) = 50
x^2 + 5x - 50 = 0
Factor:
(x-5)(x+10) = 0
x = 5 or x = -10
========================
Cheers,
Stan H.